P5.5. Eunice Hospitals preadjusted trial balance at December 31, 20x1, included the following accounts, among others: Acct.
Answer #1

1. Prepaid Insurance Dr 3840

To Insurance Account 3840

Prepaid Insurance of 5760 is for 3 years so Prepaid Insurance for 2 years will be 5760 / 3 x 2 = 3840

2.Prepaid Rent Dr 1125

To Rent Account 1125

Rent paid in advance for 1 year on 1st June 2700 so 5 months rent will be advance so 2700 / 12 x 5 = 1125

3.Prepaid Interest Dr 4000

To Interest Account 4000

Interest paid in advance for 6 months so 4800 / 6 x 5 =4000

4.Salary & Wages Account Dr 27400

To Accrued Salary & Wages Payable 27400

5.Interest on bonds Account Dr 4500

To Accrued Interest Payable 4500

100000 x 6 / 100 x 9 /12 = 4500